For what values of k does 3x^2 -kx + 2 = 0 have equal roots
Can you please help me out , thanks so much in advance:)

Answer by mananth(16946)

3x^2 -kx + 2 = 0 have equal roots

Therefore discriminant = b^2-4ac =0
b=-k, a=3,c=2
b^2-4ac= (-k)^2-4*3*2=0
k^2-24=0
k^2=24
k= +/- 2sqrt(6)